Nuclear grade Uranium is the key material used in nuclear reactors. In order to understand the property of Uranium isotopes scientists use an equipment which is called Mass Spectrometer. They injected a uranium iron beam with a speed 3.00 x 105 m/s into a 0.6 T magnetic field. Beam velocity and the magnetic field are perpendicular each other. Calculate the separation between impacts points for two isotopes which have masses 3.9508 x 10-25 kg and 3.901 x 10-25.kg. Assume the charge of uranium iron is 1.6 x 10-19 C.
Answer #1

Step 1: Find radius of circular path of isotope 1 in uniform magnetic field, using force balance on the istope 1's atoms:

Force on Uranium isotope ions will be:

Force due to circular motion = Force due to magnetic field

Fc = Fm

m1*V^2/R1 = q*V*B

R1 = m1*V/(q*B)

m1 = mass of isotope 1 = 3.9508*10^-25 kg

V = Speed of ions = 3.00*10^5 m/s

q = charge on singly-ionized atoms = +e = 1.6*10^-19 C

B = magnetic field = 0.6 T

So,

R1 = 3.9508*10^-25*3.00*10^5/(1.6*10^-19*0.6)

R1 = 1.234625 m

Step 2: Find out the radius of path followed by Isotope 2:

Since we know that:

m2*V^2/R2 = q*V*B

R2 = m2*V/(q*B)

m2 = mass of isotope 2 = 3.901*10^-25 kg

R2 = 3.901*10^-25*3.00*10^5/(1.6*10^-19*0.6)

R2 = 1.2190625 m

Step 3: Now impact point of isotope will be equal to diameter of circular path, So

So separation between impact point of both isotopes will be:

d = d2 - d1 = 2*(R1 - R2)

d = 2*(1.234625 - 1.2190625)

d = 0.031125 m = 3.1125*10^-2 m

Let me know if you've any query.